OpenStudy (anonymous):

The volume of a cone that fits exactly inside a cylinder is 3 cubic feet. What is the volume of the cylinder? (5 points) Select one: a. 6 cubic feet b. 9 cubic feet c. 12 cubic feet d. 18 cubic feet

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

the volume of cone =1/3 volume of the cylinder volume of the cylinder=3 × volume of cone =3*3=9

OpenStudy (anonymous):

thank you so answer is 9 correct

OpenStudy (ikram002p):

good luck ^^

OpenStudy (anonymous):

ok i have a couple more